OpenStudy (j.f):

The letters a, b, c, d, e, f, g, h represent consecutive integers. If 2b+f=13, what is the value of b? so the numbers are consecutive (1+2+3...). There is a space of 3 numbers between b and f. If I would say 2(3)+(6) this would equal 12. This is the closest I can get so there must be some other way to get this answer

OpenStudy (j.f):

Unless it doesn't matter that there is 3 numbers between b and f I could say 2(3)+7= 13

Parth (parthkohli):

Actually they're 4 apart.

OpenStudy (rational):

a, `b`, c, d, e, `f`, g, h c = b+1 d = b+1+1 e = b+1+1+1 f = b+1+1+1+1

OpenStudy (rational):

2b + f = 13 2b + (b+1+1+1+1) = 13 you can solve b

OpenStudy (j.f):

2(3)+7=13